OpenStudy (anonymous):

Find the lateral area for the regular pyramid. slant height = 6, length of side = 4

OpenStudy (anonymous):

anyone know how to do this?

OpenStudy (anonymous):

 if you need the area of the triangle .. sqrt(40) * 4 / 2 = sqrt(40)*2 = 4*sqrt(10)